# Memory Intelligence Agent

### 06 Apr 2026 v1

@@ 3.2.3 Experience Consolidation
An LLM Judger is utilized to evaluate the final result. The Memory Manager then compresses the current search trajectories and updates the memory units. First, to reduce storage and retrieval burdens, images are compressed into captions, and verbose trajectories are abstracted into structured workflow summaries to form the new memory, as shown in Figure 4. A Qwen3-32B LLM is utilized to complete this compression task. After that, we calculate the semantic similarity between the new memory and existing memory units to conduct a high semantic similarity knowledge replacement. If no similar memory unit is found, the new memory is stored as a new memory unit. Concurrently, the value rewards and frequency counts of relevant units are also updated. Finally, we propose a real-time exploration and update strategy to retrain the Planner with the current batch of questions, trajectories, and results, thereby internalizing episodic memory into parametric memory. This strategy enables further memory compression and efficient memory extraction. After training, the memory units will be selectively cleared. This prevents saturation while retaining critical information.
@@ 4.2 Main Result
As shown in Table 3, we find that our proposed MIA achieves the highest overall performance among open-source models, reaching an average accuracy of 53.6. Compared to the previous best memory-based method, MIA improves the average accuracy by 5.46, including specific increases of 3.3 on FVQA-test, 6.4 on the multi-hop task LiveVQA, and an impressive 9.1 on the highly challenging custom task In-house 1. Additionally, we observe a critical phenomenon: traditional contextual memory methods (e.g., RAG, Mem0, and A-Mem) generally underperform the "No Memory" baseline. This validates that long memory contexts introduce noise, leading to performance degradation. Although recent advanced memory methods (e.g., ReasoningBank, ExpeL, and Memento) abstract memory into high-level guidance to mitigate this issue, they still struggle to fully internalize historical experiences. Notably, MIA effectively bridges this gap through its dual-memory mechanism and online parameter updating, achieving an optimal balance and the highest accuracy among all memory-based approaches.
@@ 4.3 Training Analysis
In Figure 7, we present the reward function curves of the Planner and the Executor during the training stage. As the number of training steps increases, the rewards of both the Planner and the Executor show an overall upward trend, although they exhibit distinct patterns. Because the reward function can provide relatively direct and stable feedback based on the action of the Planner, the reward curve of the Planner shows a clear increase as training progresses. In contrast, the Executor’s actions are indirect through the results produced by the Planner. Consequently, the feedback provided by the reward function is also indirect and unstable, which can reduce the quality of the feedback. As a result, the reward curve of the Executor increases more slowly as training grows.
@@ 4.4 Generalization to Closed-Source Executors
To evaluate the generalizability of MIA, we replace the open-source Executor with the most powerful closed-source models (GPT-5.4 (Singh et al., 2025), Gemini-3-Flash (Google DeepMind, 2026), and Claude-Sonnet-4.6 (Anthropic, 2025) while removing TTL. As shown in Figure 8, MIA achieves consistent improvements across all three models on both LiveVQA (multimodal) and HotpotQA (text-only) benchmarks. Furthermore, the improvement margin is inversely correlated with the base capability of the Executor: GPT-5.4 benefits the most (+8.9+8.9 on LiveVQA, +6.4+6.4 on HotpotQA), followed by Gemini-3-Flash (+3.1+3.1 on LiveVQA, +2.6+2.6 on HotpotQA) and Claude-Sonnet-4-6 (+1.8+1.8 on LiveVQA, +1.7+1.7 on HotpotQA).
@@
These results demonstrate that MIA possesses exceptional generalization. The method consistently enhances performance across both open-source and closed-source models of varying scales. Additionally, our approach achieves state-of-the-art gains. It notably yields significant improvements on GPT-5.4 and delivers higher improvements on other SOTA LLMs, including Gemini-3-Flash and Claude-Sonnet-4.6.
@@ 4.5 Tool Call Analysis
Figure 9: Tool call distribution analysis. Each dot represents an individual task execution (gray indicates failures; colored indicates successful runs). The scatter, box, and half-violin plots illustrate the frequency of tool calls among successful executions.
Figure 9 presents the distribution of tool calls across tasks sampled from TTL, where colored points denote successful executions. From the Figure, we can draw the following conclusions: (1) Memory is essential: Models lacking a memory mechanism and exhibiting very low tool usage (e.g., No-Memory and RAG) achieve the poorest accuracy. This is because they are used to conduct reasoning within limited tool calls. Without memory, they cannot learn to frequently recall previous tool interactions during multi-turn reasoning processes, leading to the worst performance. (2) Planning for the current query is more effective than relying solely on historical experiences: Methods based on long-context memory (e.g., ReasoningBank and ExpeL) show weaker performance compared with approaches such as Memento and MIA, which incorporate an explicit planner. (3) Combining heterogeneous memory with continual learning during test time yields the best results: By integrating multiple memory components together with test-time learning, MIA achieves the strongest performance among all memory-based systems.

### 09 Apr 2026 v3

@@ 3.2.3 Experience Consolidation
An LLM Judger is utilized to evaluate the final result. The Manager then compresses the current trajectories and updates the memory units. First, to reduce storage and retrieval burdens, images are compressed into captions, and verbose trajectories are abstracted into structured workflow summaries to form the new memory, as shown in Figure 4. A Qwen3-32B LLM is utilized to complete this compression task. After that, we calculate the semantic similarity between the new memory and existing memory units to conduct a high semantic similarity knowledge replacement. If no similar memory unit is found, the new memory is stored as a new memory unit. Concurrently, the value rewards and frequency counts of relevant units are also updated. Finally, we propose a real-time exploration and update strategy to retrain the Planner with the current batch of questions, trajectories, and results, thereby internalizing episodic memory into parametric memory. This strategy enables further memory compression and efficient memory extraction. After training, the memory units will be selectively cleared. This prevents memory explosion while retaining critical information.
@@ 4.2 Main Result
As shown in Table 3, we find that our proposed MIA achieves the highest overall performance among open-source models, reaching an average accuracy of 53.6. Compared to the previous best memory-based method, MIA improves the average accuracy by 5.5, including specific increases of 3.3 on FVQA-test, 6.4 on the multi-hop task LiveVQA, and an impressive 9.1 on the highly challenging custom task In-house 1. Additionally, we observe a critical phenomenon: traditional contextual memory methods (e.g., RAG, Mem0, and A-Mem) generally underperform the "No Memory" baseline. This validates that long memory contexts introduce noise, leading to performance degradation. Although recent advanced memory methods (e.g., ReasoningBank, ExpeL, and Memento) abstract memory into high-level guidance to mitigate this issue, they still struggle to fully internalize historical experiences. Notably, MIA effectively bridges this gap through its dual-memory mechanism and online parameter updating, achieving an optimal balance and the highest accuracy among all memory-based approaches.
@@ 4.3 Training Analysis
In Figure 7, we present the reward function curves of the Planner and the Executor during the training stage. As the number of training steps increases, the rewards of both the Planner and the Executor show an overall upward trend, although they exhibit distinct patterns. Because the reward function can provide relatively direct and stable feedback based on the action of the Executor, the reward curve of the Executor shows a clear increase as training progresses. In contrast, the Planner’s actions are indirect through the results produced by the Executor. Consequently, the feedback provided by the reward function is also indirect and unstable, which can reduce the quality of the feedback. As a result, the reward curve of the Planner increases more slowly as training grows.
@@ 4.4 Generalization to Closed-Source Executors
To evaluate the generalizability of MIA, we replace the open-source Executor with the most powerful closed-source models (GPT-5.4 (Singh et al., 2025), Gemini-3-Flash (Google DeepMind, 2026), and Claude-Sonnet-4.6 (Anthropic, 2025). In this experimental setup, we only perform training on the Planner (as the way in the TTL stage) since the Executor’s parameters are inaccessible. Meanwhile, we also update the non-parametric memory, which supports plan generation and is continuously enriched with the Executor’s execution trajectories. As shown in Figure 8, MIA achieves consistent improvements across all three models on both LiveVQA (multimodal) and HotpotQA (text-only) benchmarks. Furthermore, the improvement margin is inversely correlated with the base capability of the Executor: GPT-5.4
@@
benefits the most (+8.9+8.9 on LiveVQA, +6.4+6.4 on HotpotQA), followed by Gemini-3-Flash (+3.1+3.1 on LiveVQA, +2.6+2.6 on HotpotQA) and Claude-Sonnet-4.6 (+1.8+1.8 on LiveVQA, +1.7+1.7 on HotpotQA). These results demonstrate that MIA possesses exceptional generalization. The method consistently enhances performance across open-source and closed-source models. Additionally, our approach achieves state-of-the-art gains. It notably yields significant improvements on GPT-5.4 and delivers higher improvements on other SOTA LLMs, including Gemini-3-Flash and Claude-Sonnet-4.6.
@@ 4.5 Tool Call Analysis
Figure 9: Tool call distribution analysis. Each dot represents an individual task execution (gray indicates failures; colored indicates successful runs). The scatter and half-violin plots illustrate the frequency of tool calls among successful executions.
Figure 9 presents the distribution of tool calls across tasks sampled from TTL, where colored points denote successful executions. From the figure, we can draw the following conclusions: (1) Memory is essential: Models lacking a memory mechanism and exhibiting very low tool usage (e.g., No-Memory) achieve the poorest accuracy. This is because models are used to conduct reasoning within limited tool calls. Without memory, they cannot effectively recall previous tool interactions during multi-turn reasoning processes, leading to the worst performance. (2) Planning for the current query is more effective than relying solely on historical experiences: Methods based on long-context memory (e.g., RAG, Mem0 and A-Mem) or meta-guidance memory (e.g., ReasoningBank and Expel) show weaker performance compared to approaches such as Memento and MIA, which incorporate an explicit planner on top of meta-guidance memory. (3) Combining heterogeneous memory with continual learning during test time yields the best results: By integrating multiple memory components with test-time learning, MIA achieves the strongest performance among all memory-based systems.